# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ IV-A Reliability Analysis
From above, we further have $\forall\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$,
$\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))\leq\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n)),$ (49)
$\mathsf{V}(\Psi(n),\mu\Psi(n))\geq\mathsf{V}(\Upsilon,\mu\Psi(n)),$ (50)
@@
Let $B(\Upsilon,\mu\Psi(n))=\frac{6\mathsf{T}\left[S_{i}(\Upsilon,\mu\Psi(n))\right]}{nV[S_{i}(\Upsilon,\mu\Psi(n))]^{\frac{3}{2}}}$ and it is easy to obtain that $B(\Upsilon,\mu\Psi(n))$ is bounded for any $\Upsilon\in[\mu^{2}\Psi,\Psi]$. Hence, for any given $10<\epsilon_{1}<$ 1, if $n$ is sufficiently large, we have
$\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}>0.$ (53)
Now, let $\gamma$ be as in (5). Obviously, $\gamma$ does not depend on $\Upsilon$. From Berry-Esseen Theorem, the inequality (IV-A) holds.
@@
$\leq$ $\Phi\left(\frac{\gamma-n\mathsf{C}(\mu,\Psi(n))-n\cdot\bar{\mathsf{C}}(\Upsilon,\Psi(n))}{\sqrt{n\mathsf{V}(\Upsilon,\Psi(n))}}\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$=$ $\Phi\left(\frac{n\cdot\mathsf{C}(\mu^{2}\Psi(n),\mu\Psi(n))-n\cdot\bar{\mathsf{C}}(\Upsilon,\Psi(n))+\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}\Phi^{-1}\big{(}\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\big{)}}{\sqrt{n\mathsf{V}(\Upsilon,\Psi(n))}}\right)+\frac{B(\Upsilon,\mu\Psi(n)}{\sqrt{n}}$
$\overset{(a)}{\leq}$ $\Phi\left(\sqrt{\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\Psi(n))}}\Phi^{-1}\big{(}\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n)}{\sqrt{n}}\big{)}\right)+\frac{B(\Upsilon,\mu\Psi(n)}{\sqrt{n}}$
$\overset{(b)}{\leq}$ $\Phi\left(\Phi^{-1}\big{(}\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n)}{\sqrt{n}}\big{)}\right)+\frac{B(\Upsilon,\mu\Psi(n)}{\sqrt{n}}$
$\leq$ $\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n)}{\sqrt{n}}+\frac{B(\Upsilon,\mu\Psi(n)}{\sqrt{n}}$
@@
In the above derivation, the step (a) follow from the facts (1): $\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))-\bar{\mathsf{C}}(\mu\Psi(n),\mu\Psi(n))<0$. (2): $c_{0}\log n>0$. (3): $\Phi$ is the cumulative distribution function. Step (b) follows from that $\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n)}{\sqrt{n}})<0$ and $\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\Psi(n))}>1$. From the arguments above, the bound is applicable with any $\Upsilon\in[\mu^{2}\Psi,\Psi]$. Due to the sphere symmetry, we have for any randomly generated codeword $\bm{x}=\bm{x}(s,w)$ from $\bar{P}_{n}^{(n)}$, $\varepsilon_{s,w}^{(1)}<\epsilon_{1}$, i.e.,
@@
Note that the bound in (55) implies $\mathbb{E}\left(\varepsilon_{w}^{(1)}\right)\leq\epsilon_{1}$.
@@
In the $\kappa\beta$ bound (Theorem 25) of [2], the codewords lie on $\mathcal{S}^{n-1}$ and are not randomly generated by any distribution. Nevertheless, the auxiliary distribution $\mathcal{N}(0,(1+P)\bm{I}_{n})$ is chosen so that normal approximation can be performed. In this work, the generating distribution $\bar{P}_{n}^{(n)}$ is chosen such that its induced output distribution $\bar{Q}^{(n)}_{m,1}$ is close to the auxiliary distribution $Q^{(n)}_{m,1}$, which makes the normal approximation accurate.
@@
In our previous work [24], the encoder adopts sequential random coding (Theorem 1 in [24]) so that the codewords are not i.i.d generated, which makes it difficult to perform resolvability analysis. In this work, the codewords are i.i.d sampled from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$, and the power of codewords ranges from $\mu^{2}n\Psi(n)$ to $n\Psi(n)$. Moreover, $\Psi(n)$ depends on $n$ to satisfy the covert constraint. As a consequence, there is an additional term $\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))$ in the normal approximation, which is different from the previous works [1][2] where the power of the codewords in these works are the same.
@@
In this case, the inequality like (56) no longer holds when $f^{*}_{\bm{Y}}$ is the pdf of the particular auxiliary distribution $Q^{(n)}_{m,1}$.
For the error probability of the second type, we have the following lemma.
@@
$\begin{split}\log M_{n}=\gamma-c_{0}\log n,\end{split}$ (58)
then
$\begin{split}\mathbb{E}\left(\varepsilon_{w}^{(2)}\right)=&\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{m,1}}\big{(}\frac{\bar{Q}^{(n)}_{m,1}}{Q^{(n)}_{m,1}}\big{)},\\ \end{split}$ (59)
$\log M_{n}=$ $nC(\mu,n)+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))$
$+$ $\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})$
@@
The conclusion is obvious from (36) and (5).
IV-B Resolvability Analysis
@@
This section presents one-shot resolvability analysis of covert communication over AWGN channels. Different from the binary input discrete channels in [23], where the resolvability analysis resorts to modified triangle inequality of divergences as quasi-metrics, the quasi-metric property of KL divergence no longer holds over the AWGN channel. The expectation of $\mathbb{D}(\hat{Q}^{(n)}_{\bm{z}}||\hat{Q}^{(n)}_{0})$ is directly decomposed into several terms with some auxiliary distributions, and each term is bounded separately.
The KL divergence $\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{n}_{0})$ is expressed as
$\begin{split}&\mathbb{D}(\hat{Q}^{(n)}_{\bm{z}}||\hat{Q}^{(n)}_{0})\\ =&\int_{z}\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{x}_{sw})\\ &\times\log\frac{\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{x}_{sw})}{MK\hat{Q}^{(n)}_{0}(\bm{z})}d\bm{z}\end{split}$ (61)
The above equation can be rewritten as
$\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{n}_{0})$
$=\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\left[\log\frac{\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})}{\hat{Q}^{(n)}_{1}(\bm{z})}+\log\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}\right.$
$+\left.\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\hat{Q}^{(n)}_{0}(\bm{z})}\right]d\bm{z}$
$=$ $\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\log\frac{\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})}{\hat{Q}^{(n)}_{1}(\bm{z})}d\bm{z}$
$+$ $\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\cdot\log\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}d\bm{z}$
$+$ $\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\hat{Q}^{(n)}_{0}(\bm{z})}d\bm{z}$ (62)
@@
Note that $\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{n}_{0})$ and the three terms in (IV-B) are all random variables as they are functions of $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$, which are randomly generated from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$. From the linearity of the expectation, we have
$\begin{split}\mathbb{E}&\left[\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{n}_{0})\right]\\ =&\mathbb{E}(\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{(n)}_{1}))+\mathbb{E}\left[\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\log\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}d\bm{z}\right]\\ +&\mathbb{E}\left[\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\hat{Q}^{(n)}_{0}(\bm{z})}d\bm{z}\right].\end{split}$ (63)
For the first term in (63), we have the following inequality from [33], which has also been used in Lemma 3 in [23] for resolvability analysis for DMCs.
$\begin{split}&\mathbb{E}(\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}||\hat{Q}^{(n)}_{1}))\\ \leq&\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{X})\bar{P}_{n}^{(n)}(\bm{X})}\left(\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\hat{Q}^{(n)}_{1}(\bm{Z})}\right)\right).\end{split}$ (64)
Define =$\mathcal{E}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{\bar{Q}^{(n)}_{1}(\bm{z})}\leq\gamma_{1}\}$. Note that in the definition of $\mathcal{E}$, we use $\bar{Q}^{(n)}_{1}$ but not $\hat{Q}^{(n)}_{1}$. For the ratio $\frac{W_{2}^{n}(\bm{z}|\bm{x})}{\bar{Q}^{(n)}_{1}(\bm{z})}$, the following lemma is useful for the follow-up analysis, the proof of which can be found in Appendix C.
@@
$\frac{W_{2}^{n}(\bm{z}|\bm{x})}{\bar{Q}^{(n)}_{1}(\bm{z})}\leq\bigg{(}\frac{\sigma^{2}+\mu\Psi(n)}{\sigma^{2}}\bigg{)}^{\frac{n}{2}}e^{\frac{n}{2\mu}}$ (65)
For the event $\mathcal{E}^{c}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{\bar{Q}^{(n)}_{1}(\bm{z})}>\gamma_{1}\}$, the following lemma provides an upper bound on its probability with proper chosen $\gamma_{1}$.
@@
$\gamma_{1}=n(1+\rho)[\mathsf{C}(\mu,\Psi,\sigma^{2})+\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})]$ (66)
@@
$\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})=\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\Psi(n)-\frac{\mu\Psi(n)}{\sigma^{2}}),$ (68)

### v5

@@ III-A Random Coding, Decoding and Reliability Analysis
From the above (50) and (51), we further have $\forall\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$, the following inequalities hold,
$\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))\leq\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))\leq\bar{\mathsf{C}}(\Psi(n),\mu\Psi(n)),$ (53)
$\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\mu\Psi(n))}=\frac{2\Psi(n)+\mu^{2}\Psi^{2}}{2\Upsilon+\mu^{2}\Psi^{2}}\geq 1,$ (54)
In addition, we have
$\bar{\mathsf{C}}(\mu\Psi(n),\mu\Psi(n))=0,$ (55)
@@
Let $B(\Upsilon,\mu\Psi(n))=\frac{6\mathsf{T}\left[S_{i}(\Upsilon,\mu\Psi(n))\right]}{V[S_{i}(\Upsilon,\mu\Psi(n))]^{\frac{3}{2}}}$ and it is easy to obtain that $B(\Upsilon,\mu\Psi(n))$ is bounded for any $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. Hence, for any given $0<\epsilon_{1}<1$, if $n$ is sufficiently large, we have
$\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}>0.$ (57)
Now, let $\gamma$ be as in (4). Obviously, $\gamma$ does not depend on $\Upsilon$. We rewrite the term $\mathbb{P}_{W_{1}^{n}(\bm{Y}|\bm{X}=\bm{x}_{0})}\left(\log\frac{W_{1}^{n}(\bm{Y}|{\bm{X}})}{{g}_{1}^{(n)}(\bm{Y})}<\gamma\right)$ in (III-A).
@@
$\overset{(a)}{\leq}$ $\Phi\left(\frac{\gamma-n\mathsf{C}(\mu,\Psi(n))-n\cdot\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))}{\sqrt{n\mathsf{V}(\Upsilon,\Psi(n))}}\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$=$ $\Phi\left(\frac{n\cdot\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))-n\cdot\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))+\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}\Phi^{-1}\big(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\big)}{\sqrt{n\mathsf{V}(\Upsilon,\Psi(n))}}\right)$
$+$ $\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$\overset{(b)}{\leq}$ $\Phi\left(\sqrt{\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\Psi(n))}}\Phi^{-1}\bigg(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\bigg)\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$\overset{(c)}{\leq}$ $\Phi\left(\Phi^{-1}\big(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\big)\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$=$ $\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
@@
In the derivation of (III-A), the step (a) follows from applying Berry-Esseen Theorem to i.i.d random variables $S_{i}(\Upsilon,\mu\Psi(n)),i=1,\cdots,n$, the step (b) follows from (53): $\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))-\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))\leq 0$, and step (c) follows from that $\Phi^{-1}\big(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\big)<0$ and (54): $\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\Psi(n))}\geq 1$. From the arguments above, the bound is applicable with any $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. Due to the sphere symmetry, we have for any randomly generated codeword $\bm{x}_{s,w}$ from $\bar{P}_{n}^{(n)}$, $\varepsilon_{s,w}^{(1)}<\epsilon_{1}$, i.e.,
@@
In the $\kappa\beta$ bound (Theorem 25) of [20], the codewords lie on the $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius $\sqrt{nP}$ and are not randomly generated by any distribution. Nevertheless, the auxiliary distribution $\mathcal{N}(0,(1+P)\bm{I}_{n})$ is chosen so that normal approximation can be performed. In this work, the generating distribution $\bar{P}_{n}^{(n)}$ is chosen such that its induced output distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ is close to the auxiliary distribution $Q^{(n)}_{\bm{Y},1}$, which makes the normal approximation accurate. The bounding of the probability of error $\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}$ relies on the fact that the $l^{2}$ norms of the codewords lie in a small interval $\mu^{2}n\Psi(n),n\Psi(n)]$. Our strategy is separately consider $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius from $\sqrt{\mu^{2}n\Psi(n)}$ to $\sqrt{n\Psi(n)}$, and for each sphere, we can utilize the sphere symmetry to simplify the analysis. The parameter $\mu$ makes the interval under control so that an universal decoding threshold $\gamma$ can be obtained. As a consequence, there is an additional term $\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))$ in the normal approximation, which is different from the previous works [24][20] where all the codewords have the same power level.
@@
In our previous work [23], the encoder adopts sequential random coding (Theorem 1 in [23]) so that the codewords are not i.i.d generated, which makes it difficult to perform resolvability analysis. In this work, the codewords are i.i.d sampled from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ that is close to Gaussian family. The resolvability analysis on bounding the induced output distribution of the codewords and the distribution of noise will depend much on this fact.
@@
In this case, the inequality like (60) no longer holds when $f^{*}_{\bm{Y}}$ is the pdf of the particular auxiliary distribution $Q^{(n)}_{\bm{Y},1}$.
Note that for the error probability of the second type $\varepsilon_{sw}^{(2)}$, we can not bound it for each codeword. Nevertheless, our strategy is divided into two steps. First, we use an existing results to bound the average of it in Lemma 6. Then, we leverage the expurgation later to discard part of codewords with large error probability of the second type. The following lemma is almost a restatement of part of Lemma 1 in [22]. The differences are: (1) The discrete memoryless channel is substituted by the AWGN channel. (2) The sum is substituted by integration. (3) We use $\log$ with base $2$ instead of base $e$.
Lemma 5.
$\mathbb{E}\left(\varepsilon_{sw}^{(2)}\right)=M\cdot 2^{-\gamma}\mathbb{E}_{P_{Y^{n}}}\bigg(\frac{P_{Y^{n}}}{Q_{Y^{n}}}\bigg).$ (62)
Applying Lemma 5 for (40) in our setting, we have
$\begin{split}\mathbb{E}\left(\epsilon_{sw}^{(2)}\right)=&M_{n}\cdot 2^{-\gamma}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{Q}^{(n)}_{\bm{Y},1}}{Q^{(n)}_{\bm{Y},1}}\bigg)\\ =&M_{n}\cdot 2^{-\gamma}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\end{split}$ (63)
The above upper bound on the average of the error probability of the second type can be further elaborated if we choose a proper parameter $M$.
@@
$\begin{split}M=M_{n}=e^{\gamma-c_{0}\log n},\end{split}$ (64)
for any positive constant $c_{0}$, then
[…]
and
$\log M=$ $nC(\mu,\Psi(n))+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))$
$+$ $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})$
@@
The proof is obvious from (63) and (4). Note that here we just provide an upper bound of the average error probability of the second type $\varepsilon_{sw}^{(2)}$. To get a reliable code, we will adopt expurgation later in Theorem 6 to construct a reliable code such that the maximum of the sum $\varepsilon_{sw}^{(1)}+\varepsilon_{sw}^{(2)}$ for all $(s,w)$ pair is small.
Remark 8.
From (65), the average of the decoding error probability of the second type depends on two elements. The first element is $\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. The integrand is the ratio of the probability density functions of the induced output distribution of the generating distribution and the reference distribution, where the latter is the induced distribution of Gaussian distribution without truncation. The second element is the size of the codebook, which has been almost determined in Lemma 4. Hence, it is necessary to control the effect of truncation to bound the ratio $\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}$, which depends on the choice of the parameter $\mu$ and has been bounded for sufficiently large $n$ in Lemma 3.
III-B Resolvability Analysis
@@
This section presents one-shot resolvability analysis of covert communication over AWGN channels, i.e., the divergence between the induced distribution of the codebook and the distribution of noise: $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. Different from the binary input discrete channels in [22], where the resolvability analysis resorts to modified triangle inequality of divergences as quasi-metrics, the quasi-metric property of KL divergence no longer holds over the AWGN channel because here the involved distributions lie on $\mathbb{R}^{n}$.
Consider the codewords $\bm{X}_{sw},s\in\langle K\rangle,w\in\langle M\rangle$ as random variables. The KL divergence $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ is expressed as
$\begin{split}&\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{z},0})\\ =&\int_{z}\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{X}_{sw})\\ &\times\log\frac{\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{X}_{sw})}{MKQ^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z},\end{split}$ (67)
which is difficult to analyze directly. We rewrite it as
$\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$
$=\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\left[\log\frac{Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}+\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\right$.
$+\left.\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}\right]d\bm{z}$
$=$ $\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$+$ $\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\cdot\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$+$ $\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$ (68)
@@
Note that the three terms in (III-B) are all random variables as they are functions of $\bm{X}_{sw},s=1,\cdots,K,w=1,\cdots,M$, which are randomly generated from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$. Our strategy of bounding $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ is as follows. We directly investigate the expectation of each term in (III-B). By showing that the expectation of each term in (III-B) is small, we prove that the probability that a set of legitimate codewords can be obtained by the random coding scheme is positive. From the linearity of the expectation, we have
$\begin{split}\mathbb{E}&\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right]\\ =&\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right]+\mathbb{E}\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\cdot\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}\right]\\ +&\mathbb{E}\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}\right].\end{split}$ (69)
The follow-up analysis will focus on upper bounding the three terms in (69), among which a large body of analysis will be taken on the first term and an upper bound of it is provided in Lemma 9. Here we outline the idea of the proof. First, we apply the following inequality (the inequality (10) in [37]), which has also been used in Lemma 3 in [22] for resolvability analysis for DMCs.
$\begin{split}&\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right]\\ \leq&\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{X})\bar{P}_{n}^{(n)}(\bm{X})}\left(\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\right).\end{split}$ (70)
Then, to upper bound the term (70), we divide the product space $\bm{X}\times\bm{Z}$ into two parts. For the first part, we show that the inner term $\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)$ for each $(\bm{X},\bm{Z})$ pair is under control. For the second part, we prove that the probability of it is exponentially decreasing. In fact, a precise upper bounding of the first part depends on the choice of the sizes of codebook and the key, which will be illustrated later in Theorem 6. Before establishing Lemma 9, some technical preparations are provided in Lemma 7 and Lemma 8. Define $\mathcal{E}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\leq\gamma_{1}\}$. Note that in the definition of $\mathcal{E}$, we use ${Q}^{(n)}_{\bm{Z},1}$ but not $\bar{Q}^{(n)}_{\bm{Z},1}$. For the ratio $\frac{W_{2}^{n}(\bm{z}|\bm{x})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}$, the following lemma is useful for the follow-up analysis, the proof of which can be found in Appendix D.
@@
$\frac{W_{2}^{n}(\bm{z}|\bm{x})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}\leq\bigg(\frac{\sigma^{2}+\mu\Psi(n)}{\sigma^{2}}\bigg)^{\frac{n}{2}}e^{\frac{n}{2\mu}}$ (71)
For the event $\mathcal{E}^{c}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$, the following lemma provides an upper bound on its probability with proper chosen $\gamma_{1}$. Similar with Lemma 4, our strategy is separately considering the codewords with equal power and choosing an universal threshold $\gamma_{1}$. As a consequence, the sphere symmetry can be leveraged to show that the probability of the event $\{\bm{z}:\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$ for each $\bm{x}$ has large-deviation bounds.
@@
$\gamma_{1}=n\mathsf{C}(\mu,\Psi,\sigma^{2})-n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})$ (72)
@@
$\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})=\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\mu\Psi(n)-\Psi(n)),$ (74)